OpenStudy (anonymous):

A GROCERY STORE SELLS FOUR DIFFERENT JARS OF PEANUT BUTTER. DETERMINE WHICH IS THE BEST BUY BASED ON THE COST PER UNIT. 16 OUNCE $2.02 20 OUNCE $2.28 26 OUNCE $2.70 38 OUNCE $3.50 I'm guessing its the 38 ounce but I'm not really sure

Vocaloid (vocaloid):

for each of the four answer choices, divide cost by units and pick the one that gives the lowest result

Vocaloid (vocaloid):

meaning: 2.01/16 = ? 2.70/26 = ? 2.28/20 = ? 3.50/38 = ? whichever of these four operations yields the lowest result is the best unit cost

OpenStudy (anonymous):

why

Vocaloid (vocaloid):

"cost per unit" means cost/units, in this case it would be price/ounces. the lowest cost per unit is the best buy

Vocaloid (vocaloid):

$2.01/16 ounces = ? $2.70/26 ounces = ? $2.28/20 ounces = ? $3.50/38 ounces = ? divide all of these out on your calculator and find the lowest result

Vocaloid (vocaloid):

or, to put it another way, the cost per unit tells us how much 1 oz of peanut butter costs for each of the 4 jars. we want to pay the least amount of money per oz. of peanut butter

OpenStudy (anonymous):

$2.01/16 ounces = 0.125625 $2.70/26 ounces = 0.103846154 $2.28/20 ounces = 0.114 $3.50/38 ounces = 0.092105263

OpenStudy (muscrat123):

so which is the cheapest?

OpenStudy (muscrat123):

(or the smallest number)

OpenStudy (anonymous):

idk

Vocaloid (vocaloid):

which of the 4 answers is the smallest? 0.125625 0.103846154 0.114 0.092105263

OpenStudy (anonymous):

38 ounces

OpenStudy (muscrat123):

yes